OpenStudy (anonymous):

Bessie took out a subsidized student loan of $5000 at a 2.4% APR, compounded monthly, to pay for her last semester of college. If she will begin paying off the loan in 10 months with monthly payments lasting for 20 years, what will be the total amount that she pays in interest on the loan?

OpenStudy (amistre64):

is 10 months a defered period?

OpenStudy (anonymous):

i think it is..., its for Personal Finance

OpenStudy (amistre64):

compound the loan amount for 10 months to determine a value to calculate payments with

OpenStudy (amistre64):

once we know the payments amount, we just add all the payments over 20 years and subtract the loan amount to determine the interest cost of the loan

OpenStudy (anonymous):

how do i do that ?

OpenStudy (anonymous):

oh wait....

OpenStudy (anonymous):



OpenStudy (anonymous):

would this be a correct part...?

OpenStudy (anonymous):

i mean... am i going in the riht direction?

OpenStudy (amistre64):

5000(1+.002) ... thats good, but we need to tack on an exponent right?

OpenStudy (anonymous):

yup... should i multiply 20*12 to make the exponent?

OpenStudy (amistre64):

lets worry about the 10 months first ... the balance to pay off is increasing during the first 10 months 5000(1.002)^10 would be the amount to payoff over 20 years

OpenStudy (anonymous):

oh ok, should i input this into my calculator?

OpenStudy (amistre64):

you can if you want .... but this only gets us to a balance that needs to be paid off; we would still need to calculate what the payment amount would be

OpenStudy (anonymous):

ah ok

OpenStudy (amistre64):

do you have a formula in your material for calculating payments?

OpenStudy (anonymous):

amorzitation?

OpenStudy (amistre64):

i was thinking annuity ....

OpenStudy (anonymous):

oh

OpenStudy (amistre64):

what is the formula for amortization? im not good with the terminology is all

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

would this be it?

OpenStudy (amistre64):

that does look like what ive seen floating about yes :) i created my own formula based off of the reoccurence of the payments ... \[B_n=B_ok^{12y}-P\frac{1-k^{12y}}{1-k}\] when Bn is zero the loan is paid off so solving for P we get \[0=B_ok^{12y}-P\frac{1-k^{12y}}{1-k}\] \[B_ok^{12y}=P\frac{1-k^{12y}}{1-k}\] \[B_ok^{12y}\frac{1-k}{1-k^{12y}}=P\]

OpenStudy (amistre64):

this simplifies even more to the formula in the book, but i like this one better

OpenStudy (anonymous):

urs? or the pic?

OpenStudy (amistre64):

so: P = (5000k^(10))*k^(12*20)(1-k)\(1-k^(12*20)), k=1+.024/12 http://www.wolframalpha.com/input/?i=%285000k%5E%2810%29%29*k%5E%2812*20%29%281-k%29%5C%281-k%5E%2812*20%29%29%2C+k%3D1%2B.024%2F12 and that is 12*20 payments to add up .... then subtract the 5000

OpenStudy (amistre64):

i perfer my own formula, i can adapt it to sooo many other things rather than trying to memorize all the different formulas they like to throw at you